OpenStudy (anonymous):

If f(x)= 2x-4, find f(1/6) My progress that I have already done. f(1/6)= 2(1/6)-4

OpenStudy (danjs):

yes

OpenStudy (danjs):

\[\frac{ 2 }{ 6 } - 4 *\frac{ 6 }{ 6 }\]

OpenStudy (danjs):

\[\frac{ 2 }{ 6 } - \frac{ 24 }{ 6 }\]

OpenStudy (danjs):

-22/6 = -11/3